# Mind the Gap: A Review of Arabic Post-Training Datasets and Their Limitations

### v1

@@ Abstract
Post-training has emerged as a crucial technique for aligning pre-trained Large Language Models (LLMs) with human instructions, significantly enhancing their performance across a wide range of tasks. Central to this process is the quality and diversity of post-training datasets. This paper presents a review of publicly available Arabic post-training datasets on the Hugging Face Hub, organized along four key dimensions: (1) LLM Capabilities (e.g., Question Answering, Translation, Reasoning, Summarization, Dialogue, Code Generation, and Function Calling); (2) Steerability (e.g., persona and system prompts); (3) Alignment (e.g., cultural, safety, ethics, and fairness), and (4) Robustness. Each dataset is rigorously evaluated based on popularity, practical adoption, recency and maintenance, documentation and annotation quality, licensing transparency, and scientific contribution. Our review revealed critical gaps in the development of Arabic post-training datasets, including limited task diversity, inconsistent or missing documentation and annotation, and low adoption across the community. Finally, the paper discusses the implications of these gaps on the progress of Arabic LLMs and applications while providing concrete recommendations for future efforts in post-training dataset development.
@@
Mind the Gap: A Review of Arabic Post-Training Datasets and Their Limitations
Mohammed Alkhowaiter1thanks: Contributed equally; contributions varied by focus.thanks: Corresponding author: mohammed@refineai.dev  Norah Alshahrani2,∗  Saied Alshahrani3,∗ Reem I. Masoud2,4,7,∗Alaa Alzahrani5,∗Deema Alnuhait6,∗ Emad A. Alghamdi8Khalid Almubarak8 1Refine AI  2ASAS AI  3University of Bisha 4University College London  5King Salman Global Academy for Arabic 6University of Illinois at Urbana-Champaign  7King Abdulaziz University 8HUMAIN
@@ 1 Introduction
Recent years there has been a growing interest in building high-quality post-training datasets to steer and enhance the capabilities of large language models (LLMs). The nature of post-training has evolved alongside advancements in AI models. Although post-training still occurs after pre-training on large text corpora, its focus has shifted. Previously, post-training often involved task-specific fine-tuning such as sentiment analysis, topic classification, or image classification with models like BERT Devlin et al. (2019). Today, it has expanded into a broader and more general concept.
@@
This shift became clear with the emergence of capabilities in LLMs, as highlighted by Brown et al. (2020), which demonstrated strong performance on various tasks through zero-shot or few-shot prompting, even without explicit task-specific training. These capabilities were further advanced by works like Ouyang et al. (2022), which aligned models to better follow user intent, enabling more engaging and coherent interactions in dialogue formats to utilize these capabilities. This trend has also extended to other languages, such as Arabic, which has witnessed significant growth through several Arabic-centric LLMs aimed at enhancing and improving post-training datasets.
@@
A variety of strategies have been utilized to develop post-training datasets tailored to Arabic-centric LLMs. For the JAIS models Sengupta et al. (2023), instruction tuning was performed using a mix of English and Arabic datasets. The Arabic portion was primarily composed of translated adaptations of widely adopted English post-training resources, including those from Wang et al. (2022), Taori et al. (2023), and Conover et al. (2023), along with template-based instruction datasets such as Muennighoff et al. (2023). In addition to these translated datasets, two original datasets—NativeQA-Ar and SafetyQA—were specifically developed to incorporate culturally and contextually relevant content for the United Arab Emirates and the wider Arab region.
@@
Huang et al. (2024) introduced an Arabic-centric LLMs, dubbed AceGPT, by continuing pre-training from Llama 2 Touvron et al. (2023). In the post-training phase, their primary focus was on localizing instructions and preference data. They generated synthetic Arabic data by prompting GPT-4 Achiam et al. (2023) directly in Arabic, which resulted in more culturally nuanced responses compared to prompts in English. Additionally, they incorporated well-known datasets such as Alpaca, Evol-Instruct, and Code-Alpaca into their Supervised Fine-tuning (SFT) mixture and generated corresponding Arabic versions using GPT-4 Achiam et al. (2023). ALLaM series of models Bari et al. (2024) were post-trained on datasets collected from public and proprietary sources, covering a diverse range of topics, including education, history, Arabic linguistics, politics, and religion. Additionally, their post-training dataset underwent multiple filtering steps to ensure high quality. A more recent methodology proposed by Fanar et al. (2025) introduced a synthetic data generation pipeline aimed at enriching post-training datasets with culturally contextualized content. Despite these significant efforts, publicly available Arabic post-training datasets remain considerably behind those of many other languages. Even the Arabic-centric models developed to date still struggle to compete closely with non-Arabic-centric models, whether open-source ones like DeepSeek and Qwen or proprietary models such as ChatGPT, Claude, and Gemini El Filali et al. (2025). A key reason behind this gap is that Arabic still underrepresented in post-training efforts Guellil et al. (2021) even though it is a native language of over 400 million speakers across 22 countries and its position as the fourth most used language on the Internet Boudad et al. (2018). This under representation is largely due to limited publication of Arabic post-training dataset. Moreover, the Arabic language has rich morphology, non-concatenative word formation, complex syntactic structures, and significant diglossia between Classical Arabic (CA), Modern Standard Arabic (MSA), and regional dialects Darwish (2014) which introduce additional layers of ambiguity. Given Arabic’s linguistic complexity, cultural richness, and global relevance Bakalla (2023); Versteegh (2014), it is essential to rethink how post-training resources are developed for the language. This paper systematically surveys existing Arabic post-training datasets, identifies key gaps in coverage and documentation, and offers recommendations to guide future dataset development that reflects the full linguistic and cultural spectrum of Arabic.
1.1 Releated Work
Dataset Many efforts have been made to develop post-training datasets, which generally follow two main methodological directions. The first involves research works that introduce new post-training recipes as part of their model development, such as in Sengupta et al. (2023), Huang et al. (2024), Liang et al. (2024), Zhu et al. (2024), Bari et al. (2024), Fanar et al. (2025), Alnumay et al. (2025) and Team (2025). These recipes typically include collecting, filtering, and creating post-training datasets, which may come in various forms either prompt-template-based, human-curated, synthetically generated, or translated from English datasets. Although these efforts incorporate post-training datasets, they offer little transparency about data sources, design choices, or alignment goals.
The second direction, though less common, focuses more exclusively on building high-quality post-training datasets without necessarily developing a new model. One example of this approach is Alyafeai et al. (2024) which centers on constructing culturally aligned Arabic post-training dataset.
Post-training Survey Most prior work on Arabic datasets has focused on pre-training corpora rather than post-training data. Zaghouani (2017) and Ahmed et al. (2022) survey general Arabic corpora, while Mashaabi et al. (2024) briefly mentions datasets used in Arabic LLMs, which are mostly proprietary and pre-training oriented. Task-specific studies like Saoudi and Gammoudi (2024) and Bensalem et al. (2024) target Question and Answering (Q&A) and toxicity detection, respectively, but do not address post-training needs.
Our work addresses these challenges by systematically reviewing Arabic post-training datasets and their limitations. Specifically, our work makes the following contributions:
@@ 1.1 Releated Work
  • We systematically reviewed datasets used for post-training and alignment of Arabic language models.
@@
  • We developed a tool to automatically extract Arabic post-training datasets from the Hugging Face Hub and evaluated each dataset across six dimensions: documentation, popularity, adoption, recency, licensing, and scientific value.
@@
  • We identified critical gaps in Arabic dataset development and offered recommendations to improve transparency, cultural relevance, and downstream usability.
@@ 2 Methodology
We exclusively collected Arabic post-training datasets’ metadata from the Hugging Face Hub, as it represents the most comprehensive and widely-adopted machine learning platform utilized by researchers, developers, and organizations worldwide. The platform’s standardized metadata structure, extensive dataset repository, and robust API support make it the optimal choice for systematic data collection. While we initially attempted to diversify our sources by including platforms such as GitHub, the number of datasets with sufficient metadata and standardized formatting was negligible compared to Hugging Face Hub’s extensive collection. Additionally, GitHub repositories often lack the structured metadata tags and consistent documentation standards essential for our automated collection methodology. Therefore, we focused exclusively on the Hugging Face Hub as our primary source to ensure data quality, consistency, and comprehensive coverage of available Arabic post-training datasets. Our dataset collection and evaluation pipeline is shown in Figure 1.
@@ 2.3 Evaluations of Datasets
We systematically evaluated Arabic post-training datasets across 12 task categories, mapped to four dimensions: (1) LLM Capabilities (e.g., Q&A, translation, reasoning and multi-step thinking, summarization, dialogue, code generation, and function calling); (2) Steerability (e.g., persona and system prompt); (3) Alignment (e.g., cultural alignment, safety, ethics, and fairness), and (4) Robustness. Each dataset was assessed using framework comprising six evaluation criteria: documentation and annotation quality, popularity, practical adoption, recency and active maintenance, licensing transparency, and scientific contribution. Each criterion utilizes a structured scoring system designed for consistency and reproducibility.
@@
Table 1: An example of the evaluation criteria and scoring system used for documentation and annotation quality. The remaining evaluation criteria and scoring rubrics are provided in Appendix A.
Evaluation Criteria Avg. Score Quality Level Documentation – Dataset card explains the usage of dataset
@@
Annotation – Metadata tags specify a task
@@ 3 Analysis and Results
We analyzed 366 datasets across 12 natural language processing (NLP) domains summarized in Table 2. Due to unbalanced group sizes and small sample sizes in certain domain categories, we present only descriptive statistics to avoid Type I and Type II errors associated with insufficient statistical power and unequal groups  Field (2017). The remainder of this section will first cover the descriptive statistics of the collected datasets, followed by the automated evaluation results for those datasets.
Figure 2: Distribution of datasets across tasks. Labels include the percentage of datasets in each task. Tasks with no datasets are shown for completeness.
@@
Figure 3: Overview of dataset quality across tasks. The subfigures present quality indicators including documentation, popularity, adoption, recency, licensing transparency, and scientific contribution. While the full taxonomy includes 12 tasks, we report results for the 9 tasks with available datasets. Persona & System Prompts, and Function Call, Code Generation, and Official Documentation are excluded as no datasets were available for these tasks.
@@ 3.1 Dataset Results
As shown in Figure 2 and detailed in Appendix B, the distribution of the datasets is highly skewed towards specific tasks. For example, translation and Q&A dominate, comprising 42.3% and 38.3% of the datasets, respectively. Summarization adds another 12.3%, while the remaining six tasks account for fewer than 30 datasets combined. Notably, Function Call, Persona Ownership, Code Generation, and Official Documentation have no datasets, revealing major gaps in current publicly available post-training resources.
3.2 Automated-Evaluation Results
We present our findings from the automated evaluation of the collected datasets, focusing on their documentation and annotation quality, popularity, practical adoption, recency maintenance, licensing transparency, and scientific contribution with detailed results shown in Appendix C.
  • Documentation Quality: Documentation standards show mixed results across tasks. Figure 3(a) demonstrates that specialized domains like Ethics, Bias, and Fairness and Robustness & Safety achieve excellent documentation quality (100% high-quality scores). However, these domains contain only 9 datasets in total, which may not adequately represent the broader landscape and could limit their applicability to diverse research contexts.
  • Popularity: Dataset popularity varies significantly across tasks. Figure 3(b) shows that traditional NLP tasks like Q&A, Translation, and Summarization include many widely-used datasets with strong community adoption. In contrast, tasks such as Dialog/Conversation and Ethics, Bias, and Fairness are dominated by low- and medium-popularity datasets, reflecting either niche applications or limited awareness in the broader community.
@@ 3.2 Automated-Evaluation Results
  • Dataset Maintenance: Maintenance practices vary considerably, highlighting inconsistent update schedules across the ecosystem. Figure 3(d) shows that newer research areas like Robustness & Safety and Ethics, Bias, and Fairness maintain current datasets, while established fields such as Summarization and Translation contain many outdated resources that lack regular maintenance cycles.
@@
  • Scientific Contribution: Research integration remains limited across the dataset landscape. Figure 3(f) indicates that most datasets lack formal scientific validation through peer-reviewed publications or DOI assignment. This gap suggests that many datasets represent individual contributions rather than systematically validated research resources, potentially affecting their quality and research impact.
Table 2: Summary of Arabic NLP Dataset Coverage and Identified Gaps
@@
Q&A Strong (140 datasets) Lacks community adoption & scientific validation
@@
4 Identified Gaps
Our analysis has identified several critical gaps that could significantly hinder Arabic NLP research and applications (see Table 2). Potential limitations include the following.
  • Limited Arabic Post-Training Data: Small coverage of Arabic post-training datasets leads to slow advancement in Arabic LLMs, and hence their applications. There are almost no Arabic datasets available for key post-training tasks such as Function Call, Persona Ownership, Code Generation, and Official Documentation. This scarcity significantly hampers the development of sophisticated Arabic language models that can perform complex tasks.
  • Poor Dataset Documentation: Poor documentation and annotation of datasets leads to invisible and inaccessible resources within the Arabic NLP community. As shown in Table 3, many valuable datasets remain uncategorized and difficult to discover, creating barriers for researchers who could benefit from existing work. This lack of proper documentation prevents the efficient reuse and building upon previous efforts in the field.
  • Lack of Dataset Maintenance: Lack of recency and maintenance characterizes the majority of Arabic datasets, with most open-source resources rarely receiving updates or maintenance for periods exceeding 12 months. This stagnation means that datasets become outdated and potentially less relevant to current research needs. The absence of regular updates also suggests a lack of sustained community support and ongoing development efforts.

### v2

# Mohammed Alkhowaiter1  Norah Alshahrani2,∗  Saied Alshahrani3,∗ Reem I. Masoud4,7,∗Alaa Alzahrani5,∗Deema Alnuhait6,∗ Emad A. Alghamdi8Khalid Almubarak8

@@ Abstract
Post-training has emerged as a crucial technique for aligning pre-trained Large Language Models (LLMs) with human instructions, significantly enhancing their performance across a wide range of tasks. Central to this process is the quality and diversity of post-training datasets. This paper presents a review of publicly available Arabic post-training datasets on the Hugging Face Hub, organized along four key dimensions: (1) LLM Capabilities (e.g., Question Answering, Translation, Reasoning, Summarization, Dialogue, Code Generation, and Function Calling); (2) Steerability (e.g., Persona and System Prompts); (3) Alignment (e.g., Cultural, Safety, Ethics, and Fairness); and (4) Robustness. Each dataset is rigorously evaluated based on popularity, practical adoption, recency and maintenance, documentation and annotation quality, licensing transparency, and scientific contribution. Our review revealed critical gaps in the development of Arabic post-training datasets, including limited task diversity, inconsistent or missing documentation and annotation, and low adoption across the community. Finally, the paper discusses the implications of these gaps on the progress of Arabic-centric LLMs and applications while providing concrete recommendations for future efforts in Arabic post-training dataset development.
@@
[Uncaptioned image]  A Review of Arabic Post-Training Datasets and Their Limitations
Mohammed Alkhowaiter1thanks: Contributed equally; contributions varied by focus.thanks: Corresponding author: mohammed@refineai.dev.   Norah Alshahrani2,∗   Saied Alshahrani3,∗ Reem I. Masoud4,7,∗Alaa Alzahrani5,∗Deema Alnuhait6,∗ Emad A. Alghamdi8Khalid Almubarak8 1Refine AI  2ASAS AI  3University of Bisha 4University College London  5King Salman Global Academy for Arabic 6University of Illinois at Urbana-Champaign  7King Abdulaziz University 8HUMAIN
@@ 1 Introduction
Recent years there has been a growing interest in building high-quality post-training datasets to steer and enhance the capabilities of Large Language Models (LLMs). The nature of post-training has evolved alongside advancements in AI models. Although post-training still occurs after pre-training on large text corpora, its focus has shifted. Previously, post-training often involved task-specific fine-tuning, such as sentiment analysis, topic classification, or image classification, with models like BERT Devlin et al. (2019). Today, it has expanded into a broader and more general concept.
@@
This shift became clear with the emergence of capabilities of LLMs, as highlighted by Brown et al. (2020), which demonstrated strong performance on various tasks through zero-shot or few-shot prompting, even without explicit task-specific training. These capabilities were further advanced by works like Ouyang et al. (2022), which aligned models to better follow user intent, enabling more engaging and coherent interactions in dialogue formats to utilize these capabilities. This trend has also extended to other languages, such as Arabic, which has witnessed significant growth through several Arabic-centric LLMs, aimed at enhancing and improving post-training datasets.
@@
A variety of strategies have been utilized to develop post-training datasets tailored to Arabic-centric LLMs. For the JAIS models Sengupta et al. (2023), instruction tuning was performed using a mix of English and Arabic datasets. The Arabic portion was primarily composed of translated adaptations of widely adopted English post-training resources, including those from Wang et al. (2022); Taori et al. (2023); Conover et al. (2023), along with template-based instruction datasets such as Muennighoff et al. (2023). In addition to these translated datasets, two original datasets—NativeQA-Ar and SafetyQA-Ar—were specifically developed to incorporate culturally and contextually relevant content for the United Arab Emirates and the wider Arab region.
@@
Huang et al. (2024) introduced an Arabic-centric LLMs, dubbed AceGPT, by continuing pre-training from Llama 2 Touvron et al. (2023). In the post-training phase, their primary focus was on localizing instructions and preference data. They generated synthetic Arabic data by prompting GPT-4 model directly in Arabic, which resulted in more culturally nuanced responses compared to prompts in English. Additionally, they incorporated well-known datasets, such as Alpaca, Evol-Instruct, and Code-Alpaca, into their Supervised Fine-tuning (SFT) mixture and generated corresponding Arabic versions using GPT-4 Achiam et al. (2023). ALLaM series of models Bari et al. (2024) were post-trained on datasets collected from public and proprietary sources, covering a diverse range of topics, including education, history, Arabic linguistics, politics, and religion. Additionally, their post-training dataset underwent multiple filtering steps to ensure high quality. A more recent methodology proposed by Fanar et al. (2025) introduced a synthetic data generation pipeline aimed at enriching post-training datasets with culturally contextualized content. Despite these significant efforts, publicly available Arabic post-training datasets remain considerably behind those of many other languages. Even the Arabic-centric LLMs developed to date still struggle to compete closely with known LLMs, whether open-source ones, like DeepSeek and Qwen, or proprietary models, like ChatGPT, Claude, and Gemini, according to the Open Arabic LLM Leaderboard by El Filali et al. (2025).
A key reason behind this gap is that Arabic still underrepresented in post-training efforts Guellil et al. (2021) even though it is a native language of over 400 million speakers across 22 countries, and its position as the fourth most used language on the Internet Boudad et al. (2018). This underrepresentation is largely due to limited publication of Arabic post-training dataset. Moreover, the Arabic language has rich morphology, non-concatenative word formation, complex syntactic structures, and significant diglossia between Classical Arabic (CA), Modern Standard Arabic (MSA), and Dialectal Arabic (DA), which introduce additional layers of ambiguity Darwish (2014). Given Arabic’s linguistic complexity, cultural richness, and global relevance Bakalla (2023); Versteegh (2014), it is essential to rethink how post-training resources are developed for the language.
This paper surveys existing Arabic post-training datasets, identifies critical gaps, addresses challenges, and offers recommendations, all to guide future Arabic post-training dataset development. We list our main contributions as the following:
@@
  • We systematically reviewed publicly open Arabic datasets used for post-training and alignment of Arabic-centric language models.
@@
  • We developed tools111www.github.com/refineaidev/mind-the-gap. to automatically extract Arabic post-training datasets from the Hugging Face Hub and evaluate each dataset across six dimensions: documentation, popularity, adoption, recency and maintenance, licensing transparency, and scientific value.
@@
  • We identified critical gaps in Arabic post-training dataset development and offered recommendations to improve transparency, cultural relevance, and downstream usability.
@@ 2 Methodology
We exclusively collected Arabic post-training datasets’ metadata from the Hugging Face Hub, as it represents the most comprehensive and widely-adopted machine learning platform utilized by researchers, developers, and organizations worldwide. While we initially attempted to diversify our sources by including platforms such as GitHub and Kaggle, the number of datasets with sufficient metadata and standardized formatting was negligible compared to Hugging Face Hub’s extensive collection. Additionally, GitHub and Kaggle datasets often lack the structured metadata tags and consistent documentation standards essential for our automated collection methodology. Therefore, we focused solely on the Hugging Face Hub as our primary source to ensure data quality, consistency, and comprehensive coverage of available Arabic post-training datasets. Our dataset collection and evaluation pipeline is shown in Figure 1.
@@ 2.3 Evaluations of Datasets
We evaluated Arabic post-training datasets across 12 task categories, mapped to four dimensions: (1) LLM Capabilities (e.g., Q&A, Translation, Reasoning and Multi-Step Thinking, Summarization, Dialogue, Code Generation, and Function Calling); (2) Steerability (e.g., Persona and System Prompt); (3) Alignment (e.g., Cultural Alignment, Safety, Ethics, and Fairness); and (4) Robustness. The selection of the 12 task categories was informed by two criteria: (1) alignment with established taxonomies in prior research, like Chen et al. (2025); Minaee et al. (2024), and (2) representation of distinct, functionally coherent areas relevant to LLM evaluation and dataset availability. Specifically, we synthesized insights from Minaee et al. (2024), who provide a broad survey of LLM capabilities across general NLP domains. This combined perspective ensured that our categories address both specialized applications, such as Code Generation, and general-purpose tasks, such as Summarization.
Each dataset was assessed using framework comprising six evaluation criteria: documentation and annotation quality, popularity, practical adoption, recency and active maintenance, licensing transparency, and scientific contribution. Each criterion utilizes a structured scoring system designed for simplicity, consistency, and reproducibility.
@@
Table 1: An detailed example of the evaluation criteria and scoring system used for evaluating documentation and annotation quality. The remaining evaluation criteria and scoring rubrics are provided in Appendix A (Table 4).
Evaluation Criteria Avg. Score Quality Level Documentation • Dataset card explains the usage of dataset
@@
Annotation • Metadata tags specify a task
@@ 3 Analysis and Results
We analyzed 366 datasets across 12 Natural Language Processing (NLP) domains, summarized in Table 3. Due to unbalanced group sizes and small sample sizes in certain domain categories, we present only descriptive statistics to avoid Type I and Type II errors associated with insufficient statistical power and unequal groups  Field (2017). The remainder of this section will first cover the descriptive statistics of the collected datasets, followed by the evaluation results for those datasets.
Figure 2: Distribution of datasets across tasks. Labels include the percentage of datasets in each task. Tasks with no datasets are shown for the sake of completeness.
@@
Figure 3: Overview of dataset quality across tasks. The subfigures present quality indicators including documentation, popularity, adoption, recency, licensing transparency, and scientific contribution. While the full taxonomy includes 12 tasks, we report results for the 9 tasks with available datasets. Persona & System Prompts, and Function Call, Code Generation, and Official Documentation are excluded as no datasets were available for those tasks.
@@ 3.1 Dataset Results
As shown in Figure 2 and detailed in Appendix B, the distribution of the datasets is highly skewed towards specific tasks. For example, Translation and Question Answering (Q&A) dominate, comprising 42.3% and 38.3% of the datasets, respectively. Summarization adds another 12.3%, while the remaining six tasks account for fewer than 30 datasets combined. Notably, Function Call, Persona Ownership, Code Generation, and Official Documentation have no datasets (zero datasets), revealing major gaps in current publicly available Arabic post-training resources.
3.2 Automated Evaluation Results
We present our findings from the automated evaluation of the collected datasets, focusing on their documentation and annotation quality, popularity, practical adoption, recency maintenance, licensing transparency, and scientific contribution, with detailed results shown in Appendix C.
  • Documentation Quality: Documentation standards show mixed results across tasks. Figure 3(a) demonstrates that specialized domains like Ethics, Bias, and Fairness and Robustness & Safety achieve excellent documentation quality (100% high-quality scores). Still, these domains contain only 9 datasets in total, which may not adequately represent the broader landscape and could limit their applicability to diverse research contexts.
  • Popularity: Dataset popularity varies significantly across tasks. Figure 3(b) shows that traditional NLP tasks, like Q&A, Translation, and Summarization, include many widely-used datasets with strong community adoption. In contrast, tasks such as Dialog/Conversation and Ethics, Bias, and Fairness are dominated by low-popularity and medium-popularity datasets, reflecting either niche applications or limited awareness in the broader community.
@@ 3.2 Automated Evaluation Results
  • Dataset Maintenance: Maintenance practices vary considerably, highlighting inconsistent update schedules across the ecosystem. Figure 3(d) shows that newer research areas like Robustness & Safety and Ethics, Bias, and Fairness maintain current datasets, while established tasks such as Summarization and Translation contain many outdated resources that lack regular maintenance cycles.
@@
  • Scientific Contribution: Research integration remains limited across the dataset landscape. Figure 3(f) indicates that most datasets lack formal scientific validation through peer-reviewed publications or DOI assignment. This gap suggests that many datasets represent individual contributions rather than systematically validated research contributions.
4 Identified Gaps
Our analysis has identified several critical gaps that could significantly hinder Arabic NLP research and applications (as per Table 3). Potential gaps and limitations include the following:
  • Limited Arabic Post-Training Data: Small coverage of Arabic post-training datasets leads to slow advancement in Arabic-centric LLMs, and hence their applications. There are almost no Arabic datasets available for key post-training tasks such as Function Call, Persona Ownership, Code Generation, and Official Documentation. Undoubtedly, this scarcity significantly hampers the development of sophisticated Arabic large language models that can perform complex tasks.
  • Poor Dataset Documentation: Poor documentation and annotation of datasets leads to invisible and inaccessible resources within the Arabic NLP community. As shown in Table 2, many valuable datasets remain uncategorized and difficult to discover, creating barriers for researchers who could benefit from existing work. This lack of proper documentation surely prevents the efficient reuse and building upon previous efforts in the field.
  • Lack of Dataset Maintenance: Lack of recency and maintenance characterizes the majority of Arabic datasets, with most open-source resources rarely receiving updates or maintenance for periods exceeding 12 months. This stagnation means that datasets become outdated and potentially less relevant to current research needs. The absence of regular updates suggests a lack of sustained community support and ongoing development efforts.
Table 3: Summary of Arabic Post-training Dataset Coverage and Key Identified Gaps
@@ 4 Identified Gaps
Question Answering (Q&A) Strong (140 datasets) Lacks community adoption & scientific validation